OpenStudy (anonymous):

How do I factor this: w^3+125

OpenStudy (rational):

Hint : \[\large 125 = 5^3\]

OpenStudy (rational):

Also recall the identity \[\large a^3+b^3=(a+b)(a^2-ab+b^2)\]

OpenStudy (rational):

does that help ? :)